# FlowLog: Efficient and Extensible Datalog via Incrementality

### v1

@@ 6. Making Datalog Robust
Datalog execution typically exhibits high sensitivity to data distribution, recursive control, join orders, etc. The volatility makes conventional optimization techniques for SQL less effective, or even inapplicable. When a suboptimal plan is chosen for a recursive rule, the iterative nature of Datalog may exacerbate the slowdown. In contrast, there has been a growing interest in techniques that makes SQL queries robust, such as sideways information passing (sip(Ives and Taylor, 2008; Zhu et al., 2017), predicate transfer (Yang et al., 2024), and diamond hardened joins (Birler et al., 2024). These approaches advocate a more pessimistic stance, prioritizing resilience against worst-case scenarios to ensure stable performance. However, robustness techniques are largely absent in Datalog systems. Integrating such techniques into FlowLog is a first step toward our vision: making Datalog execution robust. In fact, for a simple $R\bowtie S$, DD’s join operator already yields robustness as it runs a symmetric hash-join that balances both inputs (i.e. no build-probe asymmetry) (McSherry et al., 2020). For multi-way joins, the optimizer of FlowLog uses a sip-style algorithm (discuss next) to stabilize iterative execution.
sip-style Algorithms. The key idea of sip-style algorithms is to pre-filter dangling tuples before the actual joins. Dangling tuples are tuples that do not participate in the final output. The Yannakakis algorithm (Yannakakis, 1981) is a classic example. It requires the join graph to be acyclic to be able to construct a join tree. Following the post-order traversal over the join tree, it applies two sequence of semijoins to pre-filter base tables: (1) a bottom-up pass that semijoins the child atoms with the parent, followed by (2) a top-down pass that uses the reduced parent atoms to further semijoin-reduce the children. The semijoins provably prune all dangling tuples and the subsequent joins exhibit robustness against bad join orders in practice (Zhu et al., 2017; Yang et al., 2024).
@@
Consider the following program from (Liagouris and Terrovitis, 2014) that describes an inference task in medical ontologies:
@@
where the atoms u, c, and s are edbs. The listing orders here are hand-picked for performance on the given dataset (Liagouris and Terrovitis, 2014). Among them, $r_{2},r_{3},r_{6}$ dominate the runtime and they are highly susceptible to bad join orders. The plan optimization in Sec. 5 can effectively handle $r_{2}$–it avoids joining u and q upfront (due to a high cost of $6$ in our cost model). However, the optimal join orders for $r_{3}$ and $r_{6}$ are obscure (note that $p,q$ are idbs in a mutual recursion). Take $r_{3}$ for example, it has a triangular join graph and all join orders are indistinguishable under our cost model (i.e. cost of $4$) and yet the chosen order is an order of magnitude faster than, say, c(y,w,z), p(x,y), p(x,w) (we will call it the bad listing order henceforth).
@@
sip Overheads. The sip rewriting introduces new intermediate (semijoin) rules, hence new IR (s): in this example, p1, p2, p3, c4, pre-filter the input tables for r3. This inevitably incurs overheads like maintaining these extra semijoin IR (s). However, this is outweighed if most of the dangling tuples are pruned. We will show in Sec. 10.4 that sip often leaves the structural optimizer (Sec. 5) with a near-worst-case scenario (i.e. all remaining tuples join), and by design, the optimizer strives for worst case optimalities. Future optimizations include incorporating techniques such as predicate transfer (Yang et al., 2024) to further mitigate the semijoin costs using lightweight Bloom filters.
@@ 7. Subplan Sharing
McSherry et al. (McSherry et al., 2020) introduced arrangements for DD to enable efficient index sharing across concurrent queries without redundant reconstruction. They demonstrated the benefits on simple Datalog programs by manually enforcing arrangement sharing.
@@
Our approach subsumes shared arrangements (McSherry et al., 2020) and further extends it to sharing common table subexpressions (CTEs), e.g. if there is another IR requesting the first column $e.0$ from $\textsf{edge}(x,y)\bowtie\textsf{reach}(y)$, the optimizer will not re-construct, but instead directly link to the output of the lower Join-FlatMap of Figure 5, say, 0x43.
@@ 9. Extensibility
Algebraic Semantics. A first extension (and optimization) is on the algebraic types for diff. Sec. 8 encodes diff as a Boolean value, to optimize batch Datalog queries. For incremental Datalog (edbs have insertions/deletions over time), we fall back on integer arithmetic as DD normally does. Recursive aggregation arises frequently in modern data analytics (Madsen et al., 2016; Wang et al., 2022; Khamis et al., 2022; Shaikhha et al., 2024). For example, this program computes connected components (CC) of an undirected graph:
@@
The program iteratively propagates labels—discarding larger ones as smaller labels are found—until all nodes in a connected component share the same label. Semi-naïve execution can not handle this recursive MIN because prior facts may be retracted. In fact, most Datalog engines require invasive changes to support this (e.g. Souffle lacks recursive aggregation to-date). In contrast, FlowLog needs only minor glue: it reuses DD’s native incremental machinery (cf. CC, SSSP in Table 1). Similar to the Boolean specialization (Sec. 8), FlowLog implemented this by baking aggregations directly into the diff field via a monoid (Green et al., 2007). Intuitively, it is a carrier set plus certain on-top operations; Booleans for batch Datalog were $({\texttt{bool}},\vee)$, antijoins or incremental Datalog use integers $({\mathbb{Z}},+/-/\cdot)$, while CC uses labels and a MIN operator $({\mathbb{Z}},\texttt{MIN})$. Selecting a monoid—and, when needed, applying lift to cast between monoids—yields efficient, general aggregation semantics without executor redesign.
Distributed Execution. A natural extension is scaling FlowLog to distributed environments. DD operators support scale-out execution by-design: they are sharded across workers (i.e. threads or machines) to attain resource saturation. Unlike Soufflé and RecStep, which are grounded in single-machine execution because of their designs, or BigDatalog (Shkapsky et al., 2016) that carefully re-designs distributed execution for recursive settings, FlowLog, as future work, can seamlessly extend beyond single-node Datalog setups.
@@ 10. Benchmarking and Experiments
This section presents experimental results of FlowLog  [^4] Artifacts are available at https://github.com/hdz284/FlowLog (source code) and https://github.com/HarukiMoriarty/FlowLog-Reproduction (benchmarks)..
Programs and Datasets. We curate a benchmark suite that, to our knowledge, is the broadest yet for modern Datalog engines. It subsumes nearly all publicly available programs and datasets used in recent publications of RecStep, Soufflé, and DDlog (Fan et al., 2019; Zhao et al., 2021; Hu et al., 2021; Arch et al., 2022; Li et al., 2022), plus several new programs/datasets we created or harvested from popular open-source projects. The suite spans across graph analytics, business intelligence, and static program analysis; and stresses diverse recursion behaviors (multi-way joins, mutual, nonlinear, deep iterations, aggregations/antijoins, etc.): (1) Bipartite is a custom program that decides if a connected undirected graph is bipartite (i.e., two-colorable), starting from an initialized blue node:
@@
(2) Graph queries: single-source Reachability (Reach), Shortest Path (SSSP), Same Generation (SG), Transitive Closure (TC), and Connected Components (CC); (3) Program analysis: Andersen, CSPA, and CSDA (Context-sensitive Point-to and Dataflow Analysis) from RecStep’s suite (Fan et al., 2019) (so are the datasets); (3) Dyck represents Dyck-2 reachability (Li et al., 2022) and uses the two largest CFPQ instances (kernel, postgre) (Li et al., 2022) [^5] https://formallanguageconstrainedpathquerying.github.io/CFPQ_Data; (4) Galen (Example 6.1 for medical ontologies), and CRDT (conflict-free replicated data types) come from McSherry’s blog [^6] https://github.com/frankmcsherry/dynamic-datalog; (5) Polonius (alias analysis) for Rust’s borrow checker from an open-sourced project [^7] https://github.com/rust-lang/polonius; (6) DOOP (Bravenboer and Smaragdakis, 2009), a popular Java analysis framework, featureing 136 rules with complicated recursions. Similar to Soufflé and Flan’s evaluation, the datasets are sampled among the largest from the DaCapo suite (Blackburn et al., 2025); (7) DDISASM is a simplified disassembly analysis program from (Flores-Montoya and Schulte, 2020) and its datasets are synthesized from the widely used SMT solvers: CVC5 (Barbosa et al., 2022) and Z3 (de Moura and Bjørner, 2008). For all benchmarks, we use integer data type for inputs; string-valued attributes are pre-hashed to integers before Datalog execution.
@@
Competing Engines. We compare against several state-of-the-art, open-source Datalog engines: (1) Soufflé compiled mode [^8] Soufflé interpreter (Hu et al., 2021) is in general 1.5× slower and hence excluded from the paper., the most adopted Datalog engine, backed by years of development and a rich set of optimizations (Hu et al., 2021; Subotic et al., 2018; Jordan et al., 2019; Arch et al., 2022); (2) RecStep (Fan et al., 2019), which attains stronger performance and scalability over similar designs such as BigDatalog; (3) DDlog (Ryzhyk and Budiu, 2019), which shares the same DD backend as FlowLog but differs in design and lacks key optimizations (e.g. memory reduction) discussed in this paper, making it a fair baseline. We also report results from (4) DuckDB (Raasveldt and Mühleisen, 2020), and (5) Umbra (Birler et al., 2024)—two highly optimized, state-of-the-art database engines that can execute a subset of our Datalog workloads in SQL. DuckDB has very recently added USING KEY optimizations (Bamberg et al., 2025) for recursive CTEs, and Umbra implements worst-case optimal join (Freitag et al., 2020) and employs query compilation which allows advanced loop compilations for recursive queries (Sichert and Neumann, 2022). However, SQL has limited recursion syntax, and as (Herlihy et al., 2025) and Table 1 show, nearly half of our benchmarks cannot be directly executed on both databases due to unsupported mutual or nonlinear recursion. We thus exclude these cases from our evaluation.
@@
Soufflé (compiled) and DDlog require per-program compilation, whereas others (e.g. FlowLog) are interpreters. Soufflé incurs a $\sim$10s compilation for small programs and 30s for larger ones such as DOOP. DDlog exhibits much longer Rust compilation (Hu et al., 2021)—often ¿100s—even for small programs, due to heavy DD dependencies.
@@
Environment Setup. We assess all competing engines in their latest releases in several dimensions: runtime, scalability under increasing thread counts, and memory usage. All experiments ran on a CloudLab virtual machine (CloudLab, 2018) of two AMD EPYC 7543 32-core processors (64 physical cores, hyper-threading), running Ubuntu 22.04 LTS with 256 GB RAM.
@@ 10.2. CPU and Memory Usage
CPU. For both 4 and 64 threads, FlowLog sustains near-100% CPU. Its high CPU usage stems from DD’s asynchrony, where all computation stages—including iterate—remain active and concurrent throughout execution (see Sec. 5.3), allowing a continuous resource saturation. DuckDB and Umbra, having mature parallelism infrastructures, also generally reach near-peak utilization. In contrast, Soufflé employs only outer-loop parallelism in its index nested loop joins; RecStep relies on the DBMS’s internal parallelism, but suffers from cross-iteration synchronizations. They both show moderate CPU usage for batch-oriented workloads such as Bipartite (Figure 7a, 4 iter.), but degrade sharply—often ¡25%—on long-tail workloads such as Galen (Figure 7c, 32 iter.). Another inefficiency is evident in Figure 7a, where Soufflé and DDlog have long single-threaded phases for initial index construction. On large programs such as DOOP, which feature many rules of hybrid characteristics, Soufflé averages $50\%$ CPU usage. DDlog—lacking dedicated support for parallelism (Wu et al., 2022; Fan et al., 2022)—displays volatile CPU consumptions.
Memory. A main inefficiency for DDlog is its large memory footprint, due to direct usage of DD without careful memory control. In all bottom panels of Figure 7, DDlog exhibits the highest memory usage (¿20GB). Umbra also shows elevated usage, especially in long-tail analyses (e.g., Figure 7b, 720 iter.) and triggers multiple OOM failures in Table 1. This is likely because Umbra does not have DuckDB’s USING KEY optimization (Bamberg et al., 2025) that inserts new facts in-place, and instead accumulates all results in memory. In contrast, DuckDB, Soufflé and RecStep incur lower memory overheads, generally staying ¡5GB. Although FlowLog builds on DD, our optimizations (Sec. 48) substantially shrink retained states: it uses average 3.5× less memory than DDlog on these workloads (e.g. 4 v.s. 20 GB on DOOP!), yet still 2–3× more than Soufflé’s lean baseline.
@@ 10.3. Parallel Scalability
On long-tail workloads, parallelism may degrade; and overheads of thread management and data exchange may outweigh the benefits. We see this in CSDA (postgresql, 720 iterations) and Polonius (1487 iterations): Soufflé and DDlog gain almost no parallel speedup. In CSDA, DuckDB and RecStep scale slightly weaker and Umbra outscales FlowLog, but they all start from a much slower single-thread baseline than FlowLog. This is largely due to these databases’ lack of continuously maintained views, forcing them to re-compute them at every iteration (Fan et al., 2022). Umbra’s occasionally competitive scaling (e.g. CSDA) is due to its morsel-driven parallelism (Leis et al., 2014), which allows it to adaptively balance work across threads.

### v3

@@ 6. Making Datalog Robust
Datalog execution typically exhibits high sensitivity to data distribution, recursive control, join orders, etc. The volatility makes conventional optimization techniques for SQL less effective, or even inapplicable. When a suboptimal plan is chosen for a recursive rule, the iterative nature of Datalog may exacerbate the slowdown. In contrast, there has been a growing interest in techniques that makes SQL queries robust, such as sideways information passing (sip(IvesT08; LIP17), predicate transfer (PT23), and diamond hardened joins (DiamondJoin24). These approaches advocate a more pessimistic stance, prioritizing resilience against worst-case scenarios to ensure stable performance. However, robustness techniques are largely absent in Datalog systems. Integrating such techniques into FlowLog is a first step toward our vision: making Datalog execution robust. In fact, for a simple $R\bowtie S$, DD’s join operator already yields robustness as it runs a symmetric hash-join that balances both inputs (i.e. no build-probe asymmetry) (arrangement). For multi-way joins, the optimizer of FlowLog uses a sip-style algorithm (discuss next) to stabilize iterative execution.
sip-style Algorithms. The key idea of sip-style algorithms is to pre-filter dangling tuples before the actual joins. Dangling tuples are tuples that do not participate in the final output. The Yannakakis algorithm (Yannakakis81) is a classic example. It requires the join graph to be acyclic to be able to construct a join tree. Following the post-order traversal over the join tree, it applies two sequence of semijoins to pre-filter base tables: (1) a bottom-up pass that semijoins the child atoms with the parent, followed by (2) a top-down pass that uses the reduced parent atoms to further semijoin-reduce the children. The semijoins provably prune all dangling tuples and the subsequent joins exhibit robustness against bad join orders in practice (LIP17; PT23).
@@
Consider the following program from (Galen) that describes an inference task in medical ontologies:
@@
where the atoms u, c, and s are edbs. The listing orders here are hand-picked for performance on the given dataset (Galen). Among them, $r_{2},r_{3},r_{6}$ dominate the runtime and they are highly susceptible to bad join orders. The plan optimization in Sec. 5 can effectively handle $r_{2}$–it avoids joining u and q upfront (due to a high cost of $6$ in our cost model). However, the optimal join orders for $r_{3}$ and $r_{6}$ are obscure (note that $p,q$ are idbs in a mutual recursion). Take $r_{3}$ for example, it has a triangular join graph and all join orders are indistinguishable under our cost model (i.e. cost of $4$) and yet the chosen order is an order of magnitude faster than, say, c(y,w,z), p(x,y), p(x,w) (we will call it the bad listing order henceforth).
@@
sip Overheads. The sip rewriting introduces new intermediate (semijoin) rules, hence new IR (s): in this example, p1, p2, p3, c4, pre-filter the input tables for r3. This inevitably incurs overheads like maintaining these extra semijoin IR (s). However, this is outweighed if most of the dangling tuples are pruned. We will show in Sec. 10.4 that sip often leaves the structural optimizer (Sec. 5) with a near-worst-case scenario (i.e. all remaining tuples join), and by design, the optimizer strives for worst case optimalities. Future optimizations include incorporating techniques such as predicate transfer (PT23) to further mitigate the semijoin costs using lightweight Bloom filters.
@@ 7. Subplan Sharing
McSherry et al. (arrangement) introduced arrangements for DD to enable efficient index sharing across concurrent queries without redundant reconstruction. They demonstrated the benefits on simple Datalog programs by manually enforcing arrangement sharing.
@@
Our approach subsumes shared arrangements (arrangement) and further extends it to sharing common table subexpressions (CTEs), e.g. if there is another IR requesting the first column $e.0$ from $\textsf{edge}(x,y)\bowtie\textsf{reach}(y)$, the optimizer will not re-construct, but instead directly link to the output of the lower Join-FlatMap of Figure 5, say, 0x43.
@@ 9. Extensibility
Algebraic Semantics. A first extension (and optimization) is on the algebraic types for diff. Sec. 8 encodes diff as a Boolean value, to optimize batch Datalog queries. For incremental Datalog (edbs have insertions/deletions over time), we fall back on integer arithmetic as DD normally does. Recursive aggregation arises frequently in modern data analytics (Flix16; WangK0PS22; POPS22; NestedRec24). For example, this program computes connected components (CC) of an undirected graph:
@@
The program iteratively propagates labels—discarding larger ones as smaller labels are found—until all nodes in a connected component share the same label. Semi-naïve execution can not handle this recursive MIN because prior facts may be retracted. In fact, most Datalog engines require invasive changes to support this (e.g. Souffle lacks recursive aggregation to-date). In contrast, FlowLog needs only minor glue: it reuses DD’s native incremental machinery (cf. CC, SSSP in Table 1). Similar to the Boolean specialization (Sec. 8), FlowLog implemented this by baking aggregations directly into the diff field via a monoid (Semirings07). Intuitively, it is a carrier set plus certain on-top operations; Booleans for batch Datalog were $({\texttt{bool}},\vee)$, antijoins or incremental Datalog use integers $({\mathbb{Z}},+/-/\cdot)$, while CC uses labels and a MIN operator $({\mathbb{Z}},\texttt{MIN})$. Selecting a monoid—and, when needed, applying lift to cast between monoids—yields efficient, general aggregation semantics without executor redesign.
Distributed Execution. A natural extension is scaling FlowLog to distributed environments. DD operators support scale-out execution by-design: they are sharded across workers (i.e. threads or machines) to attain resource saturation. Unlike Soufflé and RecStep, which are grounded in single-machine execution because of their designs, or BigDatalog (DBLP:conf/sigmod/ShkapskyYICCZ16) that carefully re-designs distributed execution for recursive settings, FlowLog, as future work, can seamlessly extend beyond single-node Datalog setups.
@@ 10. Benchmarking and Experiments
This section presents experimental results of FlowLog  [^4] Artifacts are available at https://github.com/hdz284/FlowLog..
Programs and Datasets. We curate a benchmark suite that, to our knowledge, is the broadest yet for modern Datalog engines. It subsumes nearly all publicly available programs and datasets used in recent publications of RecStep, Soufflé, and DDlog (recstep; David2021; souffleInterp; DBLP:conf/lopstr/ArchHZSS22; dyck22), plus several new programs/datasets we created or harvested from popular open-source projects. The suite spans across graph analytics, business intelligence, and static program analysis; and stresses diverse recursion behaviors (multi-way joins, mutual, nonlinear, deep iterations, aggregations/antijoins, etc.): (1) Bipartite is a custom program that decides if a connected undirected graph is bipartite (i.e., two-colorable), starting from an initialized blue node:
@@
(2) Graph queries: single-source Reachability (Reach), Shortest Path (SSSP), Same Generation (SG), Transitive Closure (TC), and Connected Components (CC); (3) Program analysis: Andersen, CSPA, and CSDA (Context-sensitive Point-to and Dataflow Analysis) from RecStep’s suite (recstep) (so are the datasets); (3) Dyck represents Dyck-2 reachability (dyck22) and uses the two largest CFPQ instances (kernel, postgre) (dyck22) [^5] https://formallanguageconstrainedpathquerying.github.io/CFPQ_Data; (4) Galen (Example 6.1 for medical ontologies), and CRDT (conflict-free replicated data types) come from McSherry’s blog [^6] https://github.com/frankmcsherry/dynamic-datalog; (5) Polonius (alias analysis) for Rust’s borrow checker from an open-sourced project [^7] https://github.com/rust-lang/polonius; (6) DOOP (BravenboerS09), a popular Java analysis framework, featureing 136 rules with complicated recursions. Similar to Soufflé and Flan’s evaluation, the datasets are sampled among the largest from the DaCapo suite (DaCapo); (7) DDISASM is a simplified disassembly analysis program from (DDISASM) and its datasets are synthesized from the widely used SMT solvers: CVC5 (cvc5) and Z3 (Z3). For all benchmarks, we use integer data type for inputs; string-valued attributes are pre-hashed to integers before Datalog execution.
@@
Competing Engines. We compare against several state-of-the-art, open-source Datalog engines: (1) Soufflé compiled mode [^8] Soufflé interpreter (souffleInterp) is in general 1.5× slower and hence excluded from the paper., the most adopted Datalog engine, backed by years of development and a rich set of optimizations (souffleInterp; IndexSouffle; TrieSouffle; DBLP:conf/lopstr/ArchHZSS22); (2) RecStep (recstep), which attains stronger performance and scalability over similar designs such as BigDatalog; (3) DDlog (ddlog), which shares the same DD backend as FlowLog but differs in design and lacks key optimizations (e.g. memory reduction) discussed in this paper, making it a fair baseline. We also report results from (4) DuckDB (DuckDB), and (5) Umbra (DiamondJoin24)—two highly optimized, state-of-the-art database engines that can execute a subset of our Datalog workloads in SQL. DuckDB has very recently added USING KEY optimizations (UsingKEY) for recursive CTEs, and Umbra implements worst-case optimal join (UmbraWCOJ) and employs query compilation which allows advanced loop compilations for recursive queries (SichertN22). However, SQL has limited recursion syntax, and as (DBLP:journals/corr/abs-2504-02443) and Table 1 show, nearly half of our benchmarks cannot be directly executed on both databases due to unsupported mutual or nonlinear recursion. We thus exclude these cases from our evaluation.
@@
Soufflé (compiled) and DDlog require per-program compilation, whereas others (e.g. FlowLog) are interpreters. Soufflé incurs a $\sim$10s compilation for small programs and 30s for larger ones such as DOOP. DDlog exhibits much longer Rust compilation (souffleInterp)—often ¿100s—even for small programs, due to heavy DD dependencies.
@@
Environment Setup. We assess all competing engines in their latest releases in several dimensions: runtime, scalability under increasing thread counts, and memory usage. All experiments ran on a CloudLab virtual machine (cloudlab) of two AMD EPYC 7543 32-core processors (64 physical cores, hyper-threading), running Ubuntu 22.04 LTS with 256 GB RAM.
@@ 10.2. CPU and Memory Usage
CPU. For both 4 and 64 threads, FlowLog sustains near-100% CPU. Its high CPU usage stems from DD’s asynchrony, where all computation stages—including iterate—remain active and concurrent throughout execution (see Sec. 5.3), allowing a continuous resource saturation. DuckDB and Umbra, having mature parallelism infrastructures, also generally reach near-peak utilization. In contrast, Soufflé employs only outer-loop parallelism in its index nested loop joins; RecStep relies on the DBMS’s internal parallelism, but suffers from cross-iteration synchronizations. They both show moderate CPU usage for batch-oriented workloads such as Bipartite (Figure 7a, 4 iter.), but degrade sharply—often ¡25%—on long-tail workloads such as Galen (Figure 7c, 32 iter.). Another inefficiency is evident in Figure 7a, where Soufflé and DDlog have long single-threaded phases for initial index construction. On large programs such as DOOP, which feature many rules of hybrid characteristics, Soufflé averages $50\%$ CPU usage. DDlog—lacking dedicated support for parallelism (DCDatalog22; DBLP:conf/datalog/FanMK22)—displays volatile CPU consumptions.
Memory. A main inefficiency for DDlog is its large memory footprint, due to direct usage of DD without careful memory control. In all bottom panels of Figure 7, DDlog exhibits the highest memory usage (¿20GB). Umbra also shows elevated usage, especially in long-tail analyses (e.g., Figure 7b, 720 iter.) and triggers multiple OOM failures in Table 1. This is likely because Umbra does not have DuckDB’s USING KEY optimization (UsingKEY) that inserts new facts in-place, and instead accumulates all results in memory. In contrast, DuckDB, Soufflé and RecStep incur lower memory overheads, generally staying ¡5GB. Although FlowLog builds on DD, our optimizations (Sec. 48) substantially shrink retained states: it uses average 3.5× less memory than DDlog on these workloads (e.g. 4 v.s. 20 GB on DOOP!), yet still 2–3× more than Soufflé’s lean baseline.
@@ 10.3. Parallel Scalability
On long-tail workloads, parallelism may degrade; and overheads of thread management and data exchange may outweigh the benefits. We see this in CSDA (postgresql, 720 iterations) and Polonius (1487 iterations): Soufflé and DDlog gain almost no parallel speedup. In CSDA, DuckDB and RecStep scale slightly weaker and Umbra outscales FlowLog, but they all start from a much slower single-thread baseline than FlowLog. This is largely due to these databases’ lack of continuously maintained views, forcing them to re-compute them at every iteration (DBLP:conf/datalog/FanMK22). Umbra’s occasionally competitive scaling (e.g. CSDA) is due to its morsel-driven parallelism (LeisBK014), which allows it to adaptively balance work across threads.